# A classical density functional theory for solvation across length scales

### v2

@@ Abstract
A general theory of solvation should also describe the polarization field induced by charged species such as ions. This is a challenging problem beyond the scope of the present study. There are, however, reasons to be optimistic. For example, the mDFT framework already demonstrates that orientational correlations of the bulk fluid can be used to construct density functionals that describe polarization [19]; introducing the results from our work should amount to a straightforward modification of the present mDFT framework. Insights from Weeks’ local molecular field theory may also prove useful in developing new functionals [60, 61, 62, 41, 63, 64] (see Refs. 65, 41 for discussions on the relationship between cDFT and local molecular field theory), as could ideas from integral equation theories (see, e.g., Refs. 66, 67). In a recent study, Sammüller et al. adopt an entirely different approach by using deep neural networks to construct the free energy functional [68]. Whether such an approach can be used for polar fluids remains to be seen, but it shows great promise. In the context of apolar solvation, our results raise the question whether machine learning can be made easier by first separating the functional into slowly- and rapidly-varying contributions. Notwithstanding the obvious areas for future development, the results of our work demonstrate a significant step toward efficient, and fully first-principles, multiscale modeling of solvation.
@@ Methods
Here we provide a brief overview of the methods used; full details are provided in the S.I. All molecular simulations have been performed with the LAMMPS simulation package[69]. To maintain the temperature, we used the CSVR thermostat[70]. For simulations of SPC/E water [39], long-ranged electrostatic interactions were evaluated using the particle–particle particle–mesh method[71], such that the RMS error in the forces was a factor of $10^{5}$ smaller than the force between two unit charges separated by 0.1 nm[72]. The rigid geometry of the SPC/E water molecules was imposed with the RATTLE algorithm[73]. For simulations of the neural network surrogate model [56] of RPBE-D3 [54, 55], we used the n2p2 package interface[74] with LAMMPS. cDFT calculations were performed with our own bespoke code, which we have made publicly available. Minimization was performed self-consistently using Picard iteration.

### v3

@@ Introduction
A general theory of solvation should also describe the polarization field induced by charged species such as ions. This is a challenging problem beyond the scope of the present study. There are, however, reasons to be optimistic. For example, the mDFT framework already demonstrates that orientational correlations of the bulk fluid can be used to construct density functionals that describe polarization [19]; introducing the results from our work should amount to a straightforward modification of the present mDFT framework. Insights from Weeks’ local molecular field theory may also prove useful in developing new functionals [63, 64, 65, 45, 66, 67] (see Refs. 68, 45 for discussions on the relationship between cDFT and local molecular field theory), as could ideas from integral equation theories (see, e.g., Refs. 69, 70). In a recent study, Sammüller et al. adopt an entirely different approach by using deep neural networks to construct the free energy functional [71]. Whether such an approach can be used for polar fluids remains to be seen, but it shows great promise. In the context of apolar solvation, our results raise the question whether machine learning can be made easier by first separating the functional into slowly- and rapidly-varying contributions. Notwithstanding the obvious areas for future development, the results of our work demonstrate a significant step toward efficient, and fully first-principles, multiscale modeling of solvation.
@@ Methods
Here we provide a brief overview of the methods used; full details are provided in the SM. All molecular simulations have been performed with the LAMMPS simulation package[72]. To maintain the temperature, we used the CSVR thermostat[73]. For simulations of SPC/E water [43], long-ranged electrostatic interactions were evaluated using the particle–particle particle–mesh method[74], such that the RMS error in the forces was a factor of $10^{5}$ smaller than the force between two unit charges separated by 0.1 nm[75]. The rigid geometry of the SPC/E water molecules was imposed with the RATTLE algorithm[76]. For simulations of the neural network surrogate model [60] of RPBE-D3 [58, 59], we used the n2p2 package interface[77] with LAMMPS. cDFT calculations were performed with our own bespoke code, which we have made publicly available. Minimization was performed self-consistently using Picard iteration.
Supplementary Material
Supplementary material includes further details on parameterizing the functional, simulation details, further details on the numerical implementation of the functional, a comparison to molecular density functional theory, and details on evaluating the local compressibility.